OpenStudy (anonymous):

In a survey of a TriDelt chapter with 50 members, 18 were taking mathematics, 35 were taking English, and 6 were taking both. How many were not taking either of these subjects?

OpenStudy (anonymous):

do you have any idea how to do this?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

ok, 20 + 35 = 55 55 - 6 = 49 54 - 49 = ?

OpenStudy (anonymous):

there are only 50 members though

OpenStudy (anonymous):

yes but those 6 students are counted twice, that's why we subtract.

OpenStudy (anonymous):

*members

OpenStudy (anonymous):

so what is 54-49=the number of members not in subjects

OpenStudy (anonymous):

how did you get that answer. I got 29+12=41 41+6=47 which =3

OpenStudy (anonymous):

sorry I meant 18+35, ok 18 people are taking math, 35 taking English we don't know who they are, if the people are taking both, so we add it, then it says 6 people are taking both so we subtract it to avoid double people in the classes. so we have 53-6=47, so 47 people are taking classes. so 50-47=3. it is 3 so good job however you got that answer.

OpenStudy (anonymous):

do you understand how I got that answer?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yeah i understand

OpenStudy (anonymous):

yw